OpenStudy (anonymous):

2x-y=1 what is the slope of this equation?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

can u guys explain it to me please

OpenStudy (anonymous):

You have to move 2x to the other side and arrange it in the form of y=mx+b and m=slope

OpenStudy (anonymous):

for starters, the slop equation is y = m x + b

OpenStudy (anonymous):

Move 2x to the other side and change the sign -y=-2x+1 Multiply all by a negative so the y won't be negative y=2x-1 The formula is y=mx+b where "m" is the slope. So look at that spot and that's your slope

OpenStudy (anonymous):

k thanks you all =)